Find the angular momentum of a 5-kg iron ball swinging in a horizontal circle at 3 m/s at the end of a 4-m rope.
1
Expert's answer
2018-04-09T03:29:12-0400
The angular momentum
L=mvr=5×3×4=60 (kg∙m^2)/s